OpenStudy (anonymous):

red light has a wave length of 684nm. what is its frequency

OpenStudy (ghazi):

speed of light = wavelength * frequency 3*10^8= 684*10^-9* frequency

OpenStudy (ghazi):

\[3*10^8=648*10^{-9}*frequency\]

OpenStudy (anonymous):

@ghazi is very correct........... can u explain it PLZ

OpenStudy (ghazi):

@sauravshakya do you want me to explain?

OpenStudy (anonymous):

Yep........ I think he didnt understand

OpenStudy (anonymous):

i want an explanation....

OpenStudy (ghazi):

consider this example If a wave with a wavelength of 2 meters is going by at a speed of 6 meters/second, then 3 complete waves go by in 1 second. That is, the frequency is 6/2 = 3 waves/second, or 3 Hz. therefore frequency = speed/ wavelength